OpenStudy (anonymous):

Calculate the pH of 6.8 × 10-8 M HCl. Report your answer to the hundredths place.

OpenStudy (anonymous):

ph = -log[H+] hint [H+] can be found out from the concentration of HCl

OpenStudy (anonymous):

@Kryten

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

Hcl -----> H^+ + Cl^-

OpenStudy (anonymous):

So H+ = 6.8 × 10-8

OpenStudy (anonymous):

Am i correct ?

OpenStudy (anonymous):

yup and from there you see that n(HCl) = n(H+) so [H+] = [HCl]

OpenStudy (anonymous):

Nw PH = -log (6.8 × 10-8) pH = 8 - log6.8 correct??

OpenStudy (anonymous):

nope pH = - log (6,8*10^-8) or if you want pH = - log 0,000000068 ;)

OpenStudy (anonymous):

ok.....thxxx....) Sir..)

OpenStudy (anonymous):

not a problem ;)